# Measuring Progress in Dictionary Learning for Language Model Interpretability with Board Game Models

### v1

@@ 1 Introduction
Mechanistic interpretability aims to reverse engineer neural networks into human-understandable components. What, however, should these components be? Recent work has applied Sparse Autoencoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2023), a scalable unsupervised learning method inspired by sparse dictionary learning to find a disentangled representation of language model (LM) internals. However, measuring progress in training SAEs is challenging because we do not know what a gold-standard dictionary would look like, as it is difficult to anticipate which ground-truth features underlie model cognition. Prior work has either attempted to measure SAE quality in toy synthetic settings (Sharkey et al., 2023) or relied on various proxies such as sparsity, fidelity of the reconstruction, and LM-assisted autointerpretability (Bills et al., 2023).
@@ 2.1 Language models for Othello and chess
In this work, we make use of LMs trained to autoregressively predict transcripts of chess and Othello games. We emphasize that these transcripts only give lists of moves in a standard notation and do not directly expose the board state. Based on behavioral evidence (the high accuracy of the LMs for predicting legal moves) and prior studies of LM representations (Li et al., 2023a; Nanda et al., 2023; Karvonen, 2024) we infer that the LMs internally model the board state, making them a good testbed setting for studying LM representations.
@@
Othello makes a natural testbed for studying emergent internal representations since the game tree is far too large to memorize. However, the rules and state are not particularly complex. Therefore, we also consider a language model trained on chess game transcripts with identical architecture, provided by Karvonen (2024). The model again had no prior knowledge of chess and was trained from scratch on 16 million human games from the Lichess chess game database (Lichess, 2024). The input to the model is a string in the Portable Game Notation (PGN) format (e.g., “1. e4 e5 2. Nf3 ...”). The model predicts a legal move in 99.8 % of cases and, similar to Othello, it has an internal representation of the board state that can be extracted from the internal activations using a linear probe (Karvonen, 2024).
@@ 2.2 Sparse autoencoders
In this paper, we train SAEs on datasets consisting of activations extracted from the residual stream after the sixth layer for both the chess and Othello models. At these layers, linear probes trained with logistic regression were accurate for classifying a variety of properties of the game board (Karvonen, 2024; Nanda et al., 2023). For training SAEs, we employ a variety of SAE architectures and training algorithms, as detailed in Section 4.
@@ 3 Measuring autoencoder quality for chess and Othello models
Many of the features learned by our SAEs reflect uninteresting, surface-level properties of the input, such as the presence of certain tokens. However, upon inspection, we additionally find many SAE features which seem to reflect a latent model of the board state, e.g., features that reflect the presence of certain pieces on particular squares, squares that are legal to play on, and strategy-relevant properties like the presence of a pin in chess (Figure 1).
@@ 3.2 Measuring SAE quality with board state properties
We will base our board reconstruction metric around the following human-interpretable way of recovering a board state from a feature activation profile; we emphasize that different ways of recovering boards from feature activations may lead to qualitatively different results. This recovery rule is based on the assumption that interpretable SAE features tend to be high precision for some subset of BSPs, in line with Templeton et al. (2024). For example, features that classify common configurations of pieces are high precision (but not necessarily high recall) for multiple BSPs. Thus, using a batch of “train” data $\mathcal{D}_{\text{train}}$ we identify, for each SAE feature $f_{i}$, all of the BSPs $g\in\mathcal{G}$ for which $\phi_{f_{i},t}$ is a high precision (of at least $0.95$) classifier. Then for each $g\in\mathcal{G}$ our prediction rule is
@@
Let $F_{1}(\bm{\mathcal{P}}(\{f_{i}(\mathbf{x})\});\mathbf{b})$ denote the $F_{1}$-score for a given board state $\mathbf{b}$, where $\bm{\mathcal{P}}(\{f_{i}(\mathbf{x})\})=\left\{\mathcal{P}_{g}(\{f_{i}(\mathbf{x})\})\right\}_{g\in\mathcal{G}}$ represents the full predicted board (containing predictions for all 64 squares) obtained from the SAE activations. Then, the average $F_{1}$-score over all board states in the test dataset $\mathcal{D}_{\text{test}}$ can be calculated as:
$\text{Rec}(\{x_{i}\},\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{test}}}\max_{t}F_{1}(\bm{\mathcal{P}}(\{f_{i}(\mathbf{x})\});\mathbf{b}),$ (5)
@@ 4.1 Standard SAEs
where $\mathbf{f}(\mathbf{x})$ is the vector of feature activations, and $\mathbf{\hat{x}}$ is the reconstruction.
For a standard SAE, our baseline training method is as implemented in the open-source dictionary_learning repository (Marks & Mueller, 2024), optimizing the loss
@@ 4.2 Gated SAEs
The $L_{1}$ penalty used in the original training method encourages feature activations to be smaller than they would be for optimal reconstruction (Wright & Sharkey, 2024). To address this, Rajamanoharan et al. (2024) introduced a modification to the original SAE architecture that separates the selection of dictionary elements to use in a reconstruction and estimating the coefficients of these dictionary elements. This results in the following gated architecture:
$\pi_{\text{gate}}(\mathbf{x}):=W_{\text{gate}}(\mathbf{x}-\mathbf{b}_{d})+\mathbf{b}_{\text{gate}}$
${\mathbf{\tilde{f}}}(\mathbf{x}):=\mathbb{I}\left[{\pi_{\text{gate}}}(\mathbf{x})>0\right]\odot\text{ReLU}(W_{\text{mag}}(\mathbf{x}-\mathbf{b}_{d})+\mathbf{b}_{\text{mag}})$
${\hat{x}}({\mathbf{\tilde{f}}}(\mathbf{x}))=W_{d}{\mathbf{\tilde{f}}}(\mathbf{x})+\mathbf{b}_{d}$
@@
$+\|\mathbf{x}-\hat{x}_{\text{frozen}}(\text{ReLU}(\pi_{\text{gate}}(\mathbf{x})))\|_{2}^{2}\Big{]}$
@@ 4.3 $p$-Annealing
Fundamentally, an $L_{1}$ penalty has been used to induce sparsity in SAE features because it serves as a convex relaxation of the true sparsity measure, the $L_{0}$-norm. The $L_{1}$-norm is the convex hull of the $L_{0}$-norm, making it a tractable alternative for promoting sparsity (Wright & Ma, 2022). However, the proxy loss function is not the same as directly optimizing for sparsity, leading to issues such as feature shrinkage (Wright & Sharkey, 2024) and potentially less sparse learned features. Unfortunately, the $L_{0}$-norm is non-differentiable and directly minimizing it is an NP-hard problem (Natarajan, 1995; Davis et al., 1997), rendering it impractical for training.
In this work, we propose the use of nonconvex $L_{p}^{p}$-minimization, with $p<1$, as an alternative to the standard $L_{1}$ minimization in sparse autoencoders (SAEs). This approach has been successfully employed in compressive sensing and sparse recovery to achieve even sparser representations (Chartrand, 2007; Wen et al., 2015; Yang et al., 2018; Wang et al., 2011). To perform this optimization, we introduce a method called $p$-annealing for training SAEs, based on the compressive sensing technique called $p$-continuation (Zheng et al., 2017). The key idea is to start with convex $L_{1}$-minimization through setting $p=1$ and progressively decrease the value of $p$ during training, resulting in closer approximations of the true sparsity measure, $L_{0}$, as $p$ approaches 00. We define the sparsity penalty for each batch $x$ as a function of the current training step $s$:
@@
Figure 3: Comparison of the coverage and board reconstruction metrics for chess SAE quality on $\mathcal{G}_{\text{strategy}}$. The metrics represent the average coverage and board reconstruction obtained across all BSPs in $\mathcal{G}_{\text{strategy}}$. The coverage score reports the mean of maximal F1 scores over BSPs. The absolute coverage scores vary significantly between strategy BSPs, as discussed in Appendix C. The top row corresponds to coverage, and the bottom row corresponds to board reconstruction. The left column contains a scatterplot of loss recovered vs. $L_{0}$, with the color scheme corresponding to the coverage score and each point representing different hyperparameters. We differentiate between SAE training methods with shapes.
@@
Changing the value of $p$ changes the scale of the $L_{p}^{p}$ norm. Without also adapting the coefficient $\lambda$, the strength of the sparsity penalty would vary too wildly across training. Empirically, we found that keeping a constant $\lambda$ would lead to far too weak of a sparsity penalty for the larger $p$’s at the start of training, making the process worse than simply training with a constant $p$ from the beginning.
Consequently, we aim to adapt the coefficient $\lambda$ such that the strength of the sparsity penalty is not changed significantly due to $p$ updates. Formally, the update step is:
@@ 5 Results
We find that standard SAEs trained using $p$-annealing consistently perform better than those trained with a constant $L_{1}$ penalty (Equation 8), as measured by existing proxy metrics and in terms of improvement in coverage (see Figure 2a and 4a). In fact, standard SAEs trained using $p$-annealing show a coverage score that is comparable to Gated SAEs trained without $p$-annealing. However, we find cases where our coverage metric disagrees with existing metrics. In Figure 2, for example, Gated SAEs perform achieve a higher loss recovered score than Standard SAEs trained using $p$-annealing. We emphasize that the training and inference of Gated SAEs is more computationally expensive, requiring 50% more compute per forward pass compared to Standard SAEs (Rajamanoharan et al., 2024).
@@ 7 Related work
Since the nineties, dictionary learning (Elad, 2010; Dumitrescu & Irofti, 2018), sparse regression (Foucart & Rauhut, 2013), and later, sparse autoencoders (Ng et al., 2011) have been extensively studied in the machine learning and signal processing literature. The seminal work of Olshausen & Field (1996) introduced the concept of sparse coding in neuroscience (see also (Olshausen & Field, 2004), building upon the earlier concept of sparse representations (Donoho, 1992) and matching pursuit (Mallat & Zhang, 1993). Subsequently, a series of works established the theoretical and algorithmic foundations of sparse dictionary learning (Elad & Bruckstein, 2002; Hoyer, 2002; Eggert & Korner, 2004; Aharon et al., 2006; Yang et al., 2009; Jung et al., 2014; Tillmann, 2014; Arora et al., 2015; Bao et al., 2015; Blasiok & Nelson, 2016; Chalk et al., 2018). Notably, Gregor & LeCun (2010) introduced LISTA, an unrolled version of ISTA (Daubechies et al., 2004) that learns the dictionary instead of having it fixed.
In parallel, autoencoders were introduced in machine learning to automatically learn data features and perform dimensionality reduction (Hinton & Zemel, 1993; Li et al., 2023b). Inspired by sparse dictionary learning, sparse autoencoders (Ng et al., 2011; Coates et al., 2011; Coates & Ng, 2011; Makhzani & Frey, 2013; Li et al., 2016) were proposed as an unsupervised learning model to build deep sparse hierarchical models of data, assuming a certain degree of sparsity in the hidden layer activations. Later, Luo et al. (2017) generalized sparse autoencoders (SAEs) to convolutional SAEs. Although the theory for SAEs is less developed than that of dictionary learning with a fixed dictionary, some progress has been made in quantifying whether autoencoders can, indeed, do sparse coding, e.g., Arpit et al. (2016); Rangamani et al. (2018); Nguyen et al. (2019).
@@
The individual computational units of neural networks are often polysemantic, i.e., they respond to multiple seemingly unrelated inputs (Arora et al., 2018). Elhage et al. (2022) investigated this phenomenon and suggested that neural networks represent features in linear superposition, which allows them to represent more features than they have dimensions. Thus, in an internal representation of dimension $n$, a model can encode $m\gg n$ concepts as linear directions (Park et al., 2023), such that only a sparse subset of concepts are active across all inputs – a concept deeply related to the coherence of vectors (Foucart & Rauhut, 2013) and to frame theory in general (Christensen et al., 2003). To identify these concepts, Sharkey et al. (2023) used SAEs to perform dictionary learning on a one-layer transformer, identifying a large (overcomplete) basis of features. Cunningham et al. (2023) applied SAEs to language models and demonstrated that dictionary features can be used to localize and edit model behavior. Marks et al. (2024) proposed a scalable method to discover sparse feature circuits, as opposed to circuits consisting of polysemantic model components, and demonstrated that a human could change the generalization of a classifier by editing its feature circuit. Recently, Kissane et al. (2024) explored autoencoders for attention layer outputs. These works have benefited from a variety of open-source libraries for training SAEs for LLM interpretability (Marks & Mueller, 2024; Bloom & Chanin, 2024; Cooney, 2023).
Improving and evaluating sparse autoencoders.
Despite the success of SAEs at extracting human-interpretable features, they fail to perfectly reconstruct the activations (Cunningham et al., 2023). One challenge in the training of SAEs with an $L_{1}$ penalty is shrinkage (or ’feature suppression’); in addition to encouraging sparsity, an $L_{1}$ penalty encourages feature activations to be smaller than they would be otherwise. Wright & Sharkey (2024) approached this problem by fine-tuning the sparse autoencoder without a sparsity penalty. Jermyn et al. (2024) and Riggs & Brinkmann (2024) explored alternative sparsity penalties to reduce feature suppression during training. Rajamanoharan et al. (2024) introduced Gated SAEs, an architectural variation for the encoder which both addresses shrinkage and improves on the Pareto frontier of $L_{0}$ vs reconstruction error. Recently, Gao et al. (2024) systematically evaluated the scaling laws with respect to sparsity, autoencoder size, and language model size.
The goal of dictionary learning in machine learning is to produce human-interpretable features and capture the underlying model’s computations (Bricken et al., 2023). However, quantitatively measuring interpretability is difficult and often involves manual inspection. Therefore, most existing work assesses the quality of SAEs along different proxy metrics: (1) The cross-entropy loss recovered, which reflects the degree to which the original loss of the language model can be recovered when replacing activations with the autoencoder predictions. (2) The $L_{0}$-norm of feature activations $\mathbb{E}_{z\sim\mathcal{D}}\left\|h(z)\right\|_{0}$, measuring the number of activate features given an input (Ferrando et al., 2024). Makelov et al. (2024) proposed to compare SAEs against supervised feature dictionaries in a natural language setting. However, this requires a significant understanding of the model’s internal computations and is thus not scalable.

### v2

@@ 1 Introduction
Mechanistic interpretability aims to reverse engineer neural networks into human-understandable components. What, however, should these components be? Recent work has applied Sparse Autoencoders (SAEs) Bricken et al. (2023); Cunningham et al. (2023), a scalable unsupervised learning method inspired by sparse dictionary learning to find a disentangled representation of language model (LM) internals. However, measuring progress in training SAEs is challenging because we do not know what a gold-standard dictionary would look like, as it is difficult to anticipate which ground-truth features underlie model cognition. Prior work has either attempted to measure SAE quality in toy synthetic settings Sharkey et al. (2023) or relied on various proxies such as sparsity, fidelity of the reconstruction, and LM-assisted autointerpretability Bills et al. (2023).
@@ 2.1 Language models for Othello and chess
In this work, we make use of LMs trained to autoregressively predict transcripts of chess and Othello games. We emphasize that these transcripts only give lists of moves in a standard notation and do not directly expose the board state. Based on behavioral evidence (the high accuracy of the LMs for predicting legal moves) and prior studies of LM representations Li et al. (2023a); Nanda et al. (2023); Karvonen (2024) we infer that the LMs internally model the board state, making them a good testbed setting for studying LM representations.
@@
Othello makes a natural testbed for studying emergent internal representations since the game tree is far too large to memorize. However, the rules and state are not particularly complex. Therefore, we also consider a language model trained on chess game transcripts with identical architecture, provided by Karvonen (2024). The model again had no prior knowledge of chess and was trained from scratch on 16 million human games from the Lichess chess game database Lichess (2024). The input to the model is a string in the Portable Game Notation (PGN) format (e.g., “1. e4 e5 2. Nf3 ...”). The model predicts a legal move in 99.8 % of cases and, similar to Othello, it has an internal representation of the board state that can be extracted from the internal activations using a linear probe (Karvonen, 2024).
@@ 2.2 Sparse autoencoders
In this paper, we train SAEs on datasets consisting of activations extracted from the residual stream after the sixth layer for both the chess and Othello models. At these layers, linear probes trained with logistic regression were accurate for classifying a variety of properties of the game board Karvonen (2024); Nanda et al. (2023). For training SAEs, we employ a variety of SAE architectures and training algorithms, as detailed in Section 4.
@@ 3 Measuring autoencoder quality for chess and Othello models
Many of the features learned by our SAEs reflect uninteresting, surface-level properties of the input, such as the presence of certain tokens. However, upon inspection, we additionally find many SAE features which seem to reflect a latent model of the board state, e.g., features that reflect the presence of certain pieces on particular squares, squares that are legal to play on, and strategy-relevant properties like the presence of a pin in chess (Figures 1 and 6).
@@ 3.2 Measuring SAE quality with board state properties
We will base our board reconstruction metric around the following human-interpretable way of recovering a board state from a feature activation profile; we emphasize that different ways of recovering boards from feature activations may lead to qualitatively different results. This recovery rule is based on the assumption that interpretable SAE features tend to be high precision for some subset of BSPs, in line with Templeton et al. (2024). For example, features that classify common configurations of pieces are high precision (but not necessarily high recall) for multiple BSPs. We use a consistent dataset of 1000 games as our training set $\mathcal{D}_{\text{train}}$ for identifying high-precision features across all Board State Properties (BSPs). An additional, separate set of 1000 games serves as our test set $\mathcal{D}_{\text{test}}$. Using the training set $\mathcal{D}_{\text{train}}$, we identify, for each SAE feature $f_{i}$, all of the BSPs $g\in\mathcal{G}$ for which $\phi_{f_{i},t}$ is a high precision (of at least $0.95$) classifier. Then for each $g\in\mathcal{G}$ our prediction rule is
@@
Let $F_{1}(\bm{\mathcal{P}}(\{f_{i}(\mathbf{x})\});\mathbf{b})$ denote the $F_{1}$-score for a given board state $\mathbf{b}$, where $\bm{\mathcal{P}}(\{f_{i}(\mathbf{x})\})=\left\{\mathcal{P}_{g}(\{f_{i}(\mathbf{x})\})\right\}_{g\in\mathcal{G}}$ represents the full predicted board (containing predictions for all 64 squares) obtained from the SAE activations. [^2] We do not score empty squares. Thus, the reconstruction score would be zero if no high precision features are active.
Then, the average $F_{1}$-score over all board states in the test dataset $\mathcal{D}_{\text{test}}$ can be calculated as
$\text{Rec}(\{x_{i}\},\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{test}}}\max_{t}F_{1}(\bm{\mathcal{P}}(\{f_{i}(\mathbf{x})\});\mathbf{b}).$ (5)
@@ 4.1 Standard SAEs
where $\mathbf{f}(\mathbf{x})$ is the vector of feature activations, and $\mathbf{\hat{x}}$ is the reconstruction. For a standard SAE, our baseline training method is as implemented in the open-source dictionary_learning repository (Marks and Mueller, 2024), optimizing the loss
@@ 4.2 Gated SAEs
The $L_{1}$ penalty used in the original training method encourages feature activations to be smaller than they would be for optimal reconstruction Wright and Sharkey (2024). To address this, Rajamanoharan et al. (2024) introduced a modification to the original SAE architecture that separates the selection of dictionary elements to use in a reconstruction and estimating the coefficients of these dictionary elements. This results in the following gated architecture:
$\pi_{\text{gate}}(\mathbf{x}):=W_{\text{gate}}(\mathbf{x}-\mathbf{b}_{d})+\mathbf{b}_{\text{gate}}$ (9)
${\mathbf{\tilde{f}}}(\mathbf{x}):=\mathbb{I}\left[{\pi_{\text{gate}}}(\mathbf{x})>0\right]\odot\text{ReLU}(W_{\text{mag}}(\mathbf{x}-\mathbf{b}_{d})+\mathbf{b}_{\text{mag}})$ (10)
${\hat{x}}({\mathbf{\tilde{f}}}(\mathbf{x}))=W_{d}{\mathbf{\tilde{f}}}(\mathbf{x})+\mathbf{b}_{d}$ (11)
@@
$+\|\mathbf{x}-\hat{x}_{\text{frozen}}(\text{ReLU}(\pi_{\text{gate}}(\mathbf{x})))\|_{2}^{2}\Big{]}.$
@@ 4.3 $p$-Annealing
Fundamentally, an $L_{1}$ penalty has been used to induce sparsity in SAE features because it serves as a convex relaxation of the true sparsity measure, the $L_{0}$-norm. The $L_{1}$-norm is the convex hull of the $L_{0}$-norm, making it a tractable alternative for promoting sparsity Wright and Ma (2022). However, the proxy loss function is not the same as directly optimizing for sparsity, leading to issues such as feature shrinkage Wright and Sharkey (2024) and potentially less sparse learned features. Unfortunately, the $L_{0}$-norm is non-differentiable and directly minimizing it is an NP-hard problem Natarajan (1995); Davis et al. (1997), rendering it impractical for training.
In this work, we propose the use of nonconvex $L_{p}^{p}$-minimization, with $p<1$, as an alternative to the standard $L_{1}$ minimization in sparse autoencoders (SAEs). This approach has been successfully employed in compressive sensing and sparse recovery to achieve even sparser representations Chartrand (2007); Wen et al. (2015); Yang et al. (2018); Wang et al. (2011). To perform this optimization, we introduce a method called $p$-annealing for training SAEs, based on the compressive sensing technique called $p$-continuation Zheng et al. (2017). The key idea is to start with convex $L_{1}$-minimization through setting $p=1$ and progressively decrease the value of $p$ during training, resulting in closer approximations of the true sparsity measure, $L_{0}$, as $p$ approaches 00. We define the sparsity penalty for each batch $x$ as a function of the current training step $s$:
@@
Figure 3: Comparison of the coverage and board reconstruction metrics for chess SAE quality on $\mathcal{G}_{\text{strategy}}$. The metrics represent the average coverage and board reconstruction obtained across all BSPs in $\mathcal{G}_{\text{strategy}}$. The coverage score reports the mean of maximal F1 scores over BSPs. The absolute coverage scores vary significantly between strategy BSPs, as discussed in Appendix D. The top row corresponds to coverage, and the bottom row corresponds to board reconstruction. The left column contains a scatterplot of loss recovered vs. $L_{0}$, with the color scheme corresponding to the coverage score and each point representing different hyperparameters. We differentiate between SAE training methods with shapes.
@@
Changing the value of $p$ changes the scale of the $L_{p}^{p}$ norm. Without also adapting the coefficient $\lambda$, the strength of the sparsity penalty would vary too wildly across training. Empirically, we found that keeping a constant $\lambda$ would lead to far too weak of a sparsity penalty for the larger $p$’s at the start of training, making the process worse than simply training with a constant $p$ from the beginning. Consequently, we aim to adapt the coefficient $\lambda$ such that the strength of the sparsity penalty is not changed significantly due to $p$ updates. Formally, the update step is:
@@ 5 Results
We find that standard SAEs trained using $p$-annealing consistently perform better than those trained with a constant $L_{1}$ penalty (Equation 8), as measured by existing proxy metrics and in terms of improvement in coverage (see Figure 2a and 4a). In fact, standard SAEs trained using $p$-annealing show a coverage score that is comparable to Gated SAEs trained without $p$-annealing. Further, we find that both $p$-annealing and Gated SAEs significantly outperform Standard SAEs in addressing the shrinkage problem Wright and Sharkey (2024), as detailed in Appendix E. However, we find cases where our coverage metric disagrees with existing metrics. In Figure 2, for example, Gated SAEs perform achieve a higher loss recovered score than Standard SAEs trained using $p$-annealing. We emphasize that the training and inference of Gated SAEs is more computationally expensive, requiring 50% more compute per forward pass compared to Standard SAEs Rajamanoharan et al. (2024).
@@ 7 Related work
Since the nineties, dictionary learning (Elad, 2010; Dumitrescu and Irofti, 2018), sparse regression (Foucart and Rauhut, 2013), and later, sparse autoencoders (Ng et al., 2011) have been extensively studied in the machine learning and signal processing literature. The seminal work of Olshausen and Field (1996) introduced the concept of sparse coding in neuroscience (see also Olshausen and Field (2004), building upon the earlier concept of sparse representations (Donoho, 1992) and matching pursuit Mallat and Zhang (1993). Subsequently, a series of works established the theoretical and algorithmic foundations of sparse dictionary learning (Elad and Bruckstein, 2002; Hoyer, 2002; Eggert and Korner, 2004; Aharon et al., 2006; Yang et al., 2009; Jung et al., 2014; Tillmann, 2014; Arora et al., 2015; Bao et al., 2015; Blasiok and Nelson, 2016; Chalk et al., 2018). Notably, Gregor and LeCun (2010) introduced LISTA, an unrolled version of ISTA (Daubechies et al., 2004) that learns the dictionary instead of having it fixed.
In parallel, autoencoders were introduced in machine learning to automatically learn data features and perform dimensionality reduction (Hinton and Zemel, 1993; Li et al., 2023b). Inspired by sparse dictionary learning, sparse autoencoders (Ng et al., 2011; Coates et al., 2011; Coates and Ng, 2011; Makhzani and Frey, 2013; Li et al., 2016) were proposed as an unsupervised learning model to build deep sparse hierarchical models of data, assuming a certain degree of sparsity in the hidden layer activations. Later, Luo et al. (2017) generalized sparse autoencoders (SAEs) to convolutional SAEs. Although the theory for SAEs is less developed than that of dictionary learning with a fixed dictionary, some progress has been made in quantifying whether autoencoders can, indeed, do sparse coding, e.g., Arpit et al. (2016); Rangamani et al. (2018); Nguyen et al. (2019).
@@
The individual computational units of neural networks are often polysemantic, i.e., they respond to multiple seemingly unrelated inputs (Arora et al., 2018). Elhage et al. (2022) investigated this phenomenon and suggested that neural networks represent features in linear superposition, which allows them to represent more features than they have dimensions. Thus, in an internal representation of dimension $n$, a model can encode $m\gg n$ concepts as linear directions Park et al. (2023), such that only a sparse subset of concepts are active across all inputs – a concept deeply related to the coherence of vectors Foucart and Rauhut (2013) and to frame theory in general Christensen et al. (2003). To identify these concepts, Sharkey et al. (2023) used SAEs to perform dictionary learning on a one-layer transformer, identifying a large (overcomplete) basis of features. Cunningham et al. (2023) applied SAEs to language models and demonstrated that dictionary features can be used to localize and edit model behavior. Marks et al. (2024) proposed a scalable method to discover sparse feature circuits, as opposed to circuits consisting of polysemantic model components, and demonstrated that a human could change the generalization of a classifier by editing its feature circuit. Recently, Kissane et al. (2024) explored autoencoders for attention layer outputs. These works have benefited from a variety of open-source libraries for training SAEs for LLM interpretability (Marks and Mueller, 2024; Bloom and Chanin, 2024; Cooney, 2023).